Fifty dash fiveFifty-five percent of households say they would feel secure if they had​ $50,000 in savings. You randomly select 88 households and ask them if they would feel secure if they had​ $50,000 in savings. Find the probability that the number that say they would feel secure is​ (a) exactly​ five, (b) more than​ five, and​ (c) at most five.
Answer #1

Population proportion p = 0.55

Sampling distribution of Sampling proportion R is normal with mean = 0.55 and std.deviation = ( (0.55*(1-0.55))/88 ) ^ 0.5 = 0.053

(a) P ( R = 5/88 ) = 0 because point probabilities in continuous space is zero.

(b) P ( R > (5/88) ) = P ( Z > -9.3 ) = 1 ( approx)

(c) P ( R <= (5/88) ) = P ( Z <= -9.3 ) = 0 (approx )